### v2

# Learning Actionable Counterfactual Explanations in Large State Spaces

Abstract
Recourse generators provide actionable insights, often through feature-based counterfactual explanations (CFEs), to help negatively classified individuals understand how to adjust their input features to achieve a positive classification. These feature-based CFEs, which we refer to as low-level CFEs, are overly specific (e.g., coding experience: $4\to 5+$ years) and often recommended in feature space that doesn’t straightforwardly align with real-world actions. To bridge this gap, we introduce three novel recourse types grounded in real-world actions: high-level continuous (hl-continuous), high-level discrete (hl-discrete), and high-level ID (hl-id) CFEs.
We formulate single-agent CFE generation methods, where we model the hl-discrete CFE as a solution to a weighted set cover problem and the hl-continuous CFE as a solution to an integer linear program. Since these methods require costly optimization per agent, we propose data-driven CFE generation approaches that, given instances of agents and their optimal CFEs, learn a CFE generator that quickly provides optimal CFEs for new agents. This approach, also viewed as one of learning an optimal policy in a family of large but deterministic MDPs, considers several problem formulations, including formulations in which the actions and their effects are unknown, and therefore addresses informational and computational challenges.
Through extensive empirical evaluation using publicly available healthcare datasets (BRFSS, Foods, and NHANES), we compare the proposed forms of recourse to low-level CFEs and assess the effectiveness of our data-driven approaches. Empirical results show that the proposed data-driven CFE generators are accurate and resource-efficient, and the proposed forms of recourse have various advantages over the low-level CFEs.
1 Introduction
Machine learning models are increasingly being used to guide high-stakes decision-making processes. Given the potential impact on individuals’ livelihoods, society demands transparency and the right to an explanation, as outlined in Articles 13–15 of the EuropeanParliament2016a General Data Protection Regulation. A critical aspect of this transparency is understanding how individuals (agents) can modify their input features to achieve a desired outcome, such as a positive label in a binary classification setting. Recourse generators that provide actionable insights offer one such solution (Wachter2017; Dandl2020; Mothilal20; Ustun19; Karimi21; Shalmali19; Karimi22).
A popular form of recourse, actionable recourse Ustun19, generates feature-based counterfactual explanations (CFEs), specifying precise adjustments to features to ensure that the new features collectively result in a positive classification. For comparison with our work, we refer to these as low-level CFEs. While helpful, low-level CFEs are overly specific and might be challenging to translate into real-world-like actions. To address this limitation, we introduce three novel forms of recourse that align with real-world actions: high-level continuous (hl-continuous), high-level discrete (hl-discrete), and high-level ID (hl-id) CFEs.
Calcium (mg): $\mathbf{309}$
Carbohydrate (gm): $\mathbf{109.45}$
Copper (mg): $\mathbf{0.425}$
Dietary fiber (gm): $\mathbf{4.1}$
Iron (mg): $\mathbf{4.08}$
Magnesium (mg): $\mathbf{96}$
$\vdots$
PhysActivity: $\mathbf{1}$
Fruits: $\mathbf{0}$
Veggies: $\mathbf{0}$
AnyHealthcare: $\mathbf{1}$
LowBP: $\mathbf{0}$
NoSmoke: $\mathbf{1}$
LowChol: $\mathbf{0}$
HealthBMI: $\mathbf{0}$
$\vdots$
Education: BSc in computer science
Coding experience: $\mathbf{4}$ years
Leadership: none
Published research: none
Calcium (mg): $\mathbf{309\to 113}$
Carbohydrate (gm): $\mathbf{109.45\to 43.376000000000005}$
Copper (mg): $\mathbf{0.425\to 0.21295000000000001}$
Dietary fiber (gm): $\mathbf{4.1\to 50.113950000000024}$
Iron (mg): $\mathbf{4.08\to 42.729460000000002}$
Magnesium (mg): $\mathbf{96\to 57}$
$\vdots$
Fruits: $\mathbf{0\to 1}$
Veggies: $\mathbf{0\to 1}$
LowBP: $\mathbf{0\to 1}$
LowChol: $\mathbf{0\to 1}$
HealthBMI: $\mathbf{0\to 1}$
$\vdots$
Coding experience: $\mathbf{4\to 5+}$ years
Leadership: $\mathbf{none\to 1+}$ years
Published research: $\mathbf{none\to 1+}$
Take “leavening agents: cream of tartar” ($ca_{1}$)
Take “fish, tuna, light, canned in water, drained solids” ($ca_{2}$)
Pickup the “lisinopril and atorvastatin prescription from the pharmacy” ($da_{1}$)
Adopt “a keto diet” ($da_{2}$)
Complete the Google AI residency program The agent profile (app1)The agent profile (app2)The agent profile (app3)The low-level CFEThe low-level CFEThe low-level CFEThe hl-continuous CFEThe hl-discrete CFEThe hl-id CFE
Figure 1: A comparative analysis of low-level CFEs with three high-level types: hl-continuous, hl-discrete, and hl-id, offering actionable insights to help negatively classified agents (with initial profiles/states in grey) achieve positive outcomes from binary classifiers: dietary changes to achieve a healthy waist-to-hip ratio (app1), guide an agent (app2) to meet wellness check criteria, and help an agent (app3) qualify for an AI junior research engineer role. In all cases, the low-level CFE precisely specifies which features to modify and by how much. The hl-continuous CFE adjusts multiple features numerically, e.g., corresponding to features in the profile (app3), $ca_{1}=[8.0,61.5,0.195,0.2,3.72,2.0,\cdots],ca_{2}=[17.0,0.0,0.05,0.0,1.63,23.0,\cdots]$ but the agent does not need to understand the exact changes to follow the CFE. The hl-discrete CFE ensures features meet specific eligibility thresholds without the agent needing to know the exact adjustments to take the CFE, e.g., $da_{1}=[0,0,0,0,1,0,1,0,\cdots],da_{2}=[0,1,1,0,0,0,0,1,\cdots]$. The hl-id CFE provides a single overarching high-level action that favorably modifies all the features it should. This figure illustrates how different types of CFEs might impact the agent’s ability to interpret and act on given recourse.
Figure 1 presents an illustrative example of hypothetical recourse across three binary classification tasks, highlighting the distinctions between low-level and high-level CFEs. Specifically, the hl-continuous CFEs involve general and predefined real-world-like actions, simultaneously modifying multiple decision-making features through numerical adjustments. Similarly, hl-discrete CFEs also stem from real-world actions and might modify several features simultaneously. However, unlike hl-continuous CFEs, which adjust feature values, hl-discrete CFEs modify feature eligibility. The hl-discrete CFEs use binary vector actions to ensure all features meet a predefined threshold, reducing decisions to unit simple yes/no questions. It is particularly efficient in scenarios where feature satisfiability depends on a set threshold, such as level-one decision-making in wellness checks (see Figure 1). Lastly, the hl-id CFE encapsulates either hl-continuous or hl-discrete CFEs into a single overarching action, encompassing all necessary changes without detailing specific feature modifications. This abstraction often conveys significant implicit information.
Our contributions. First, we introduce three forms of recourse: hl-continuous, hl-discrete, and hl-id CFEs, that bridge the gap between feature-based and real-world actions (Sections 1 and 2). Second, we propose single-agent CFE generation methods that leverage predefined, real-world-like actions to generate optimal CFEs. Specifically, we formulate single-agent hl-discrete CFE generation as a weighted set cover problem and single-agent hl-continuous CFE generation as an integer linear programming (ILP) problem (Section 3).
Third, we propose data-driven approaches that, given instances of agents and their optimal CFEs (agent–CFE dataset), learn a CFE generator that will quickly provide optimal hl-continuous, hl-discrete, and hl-id CFEs for new agents. Unlike the expensive single-agent CFE generation approach, the data-driven methods are more computationally friendly. Additionally, these methods are especially favorable when historical instances of agents and their optimal CFE data are available or aggregatable, recourse generators can operate independently of decision-makers, query access to the classifier is restricted or unavailable, and the actions along with their costs and explicit effects on features are unknown (Section LABEL:sec:data_driven_gen). To the best of our knowledge, these alternative forms of recourse and the data-driven approach for generating CFEs are novel contributions.
Finally, we conduct extensive experiments on $11+$ agent–CFE datasets derived from real-world healthcare datasets: BRFSS, Foods, and NHANES. We chose these datasets due to the availability of a wealth of publicly accessible data, which allows us to effectively demonstrate the benefits of incorporating real-world-like actions in CFE generation and sufficiently explore the impact of a larger action space (or action grid according to Ustun19), enabled by the high number of actionable features with broad value ranges. Alongside these real-world datasets, we also include $11+$ fully-synthetic agent–CFE datasets for comparison. We extensively compare the low-level CFEs with the hl-continuous and hl-discrete CFEs, and provide an in-depth analysis of the performance of the data-driven hl-continuous, hl-discrete, and hl-id CFE generators under various settings (Sections LABEL:sec:exp_setup and LABEL:sec:results). Our code can be accessed here: link.
2 Background
We consider a binary classification setting, where an agent with state $\mathbf{x}$ receives either a positive (desirable) or negative (undesirable) classification under a model $f(\mathbf{x})$. Although we focus on this setting, our proposed data-driven CFE generation framework generalizes to other scenarios. Given an agent state $\mathbf{x}$ with an undesirable model outcome, the objective of the CFE generator is to provide the agent with information that they can act on to achieve a desirable classification under the model. Actionable recourse (Ustun19), generates low-level CFEs where each action in the CFE precisely specifies how much the agent should add or subtract from a specific feature to ensure that, collectively, the new features (state) result in the agent receiving a desirable model outcome.
The low-level CFE generator. Ustun19 proposed an ILP-based low-level CFE generator (Equation 1) that generates a low-level CFE to help an agent change an undesirable model outcome to a desirable one.
$\min$ $\text{cost}(\mathbf{a};\mathbf{x})$ (1)
s.t. $f(\mathbf{x}+\mathbf{a})=\hat{y}^{\star}$
$\mathbf{a}\in A(\mathbf{x}),$
where $\hat{y}^{\star}$ is the desired model outcome, $A(\mathbf{x})$ denotes the set of feasible actions given the input $\mathbf{x}$, and the function $\text{cost}(\cdot;\mathbf{x}):A(\mathbf{x})\to\mathbb{R}_{+}$ encodes the preferences between these actions. When Equation 1 is feasible, the optimal actions that modify the features (i.e., $\mathbf{x}+\mathbf{a}$) and lead to a desirable model outcome are recommended to the agent (cf. Figure 1). We refer the reader to Ustun19 for a more detailed description and to Appendix LABEL:subsec:app_lowlevel_cfe for dataset-specific experimental setup and supplementary examples of low-level CFEs.
Shortcomings of low-level CFEs and their generators. We note two limitations of low-level CFEs in comparison to our proposed forms of recourse (hl-continuous, hl-discrete, and hl-id CFE), and two, as we compare the low-level CFE generator to the proposed data-driven CFE generation approach (Verma2020CounterfactualEF; Karimi22; hiddenSolon).
First, low-level CFEs are feature-based and highly specific (e.g., in Figure 1, app1, carbohydrate (gm): $109.45\to 43.376000000000005$), which may overwhelm agents and introduce additional costs to translate the CFE into implementable steps. In contrast, our proposed CFEs are better aligned with real-world scenarios, offering what you see is what you get actionable insights (see Figure 1). Furthermore, with low-level CFEs, details about the actions the agent implements (the number of them needed, which features they would simultaneously modify, and costs to incur) are often unknown beforehand, potentially leading to a misleading price of recourse and related metrics such as sparsity (few modified features) and proximity (closeness of final state to initial state) (hiddenSolon).
Second, although a CFE (e.g., complete the Google AI residency program in Figure 1) could have been optimal for several agents with different but close profiles (e.g., one has coding experience: $4$ and another $3$), the low-level CFE being too specific, would give the agents different CFEs (coding experience: $4\to 5+$ years and coding experience: $3\to 5+$ years). In contrast, our proposed recourse methods are both agent-specific (tailored to an agent’s initial state) and generalizable (providing similar recommendations to agents with comparable profiles).
Lastly, low-level CFE generators operate on a single-agent basis (Verma2020CounterfactualEF; Karimi22), making them relatively expensive. In contrast, once trained, our proposed data-driven approach can generate CFEs for multiple new agents without the need for generator re-optimization. Moreover, low-level CFE generation becomes very limited, sometimes impractical, in scenarios where access to critical information, such as sufficient training data, classifier parameters, or a comprehensive list of actions and their costs, is restricted or inaccessible. Our proposed data-driven CFE generation would work better in this case, provided there is access to historical mappings of agents and their optimal CFEs, e.g., healthcare interventional data or high school counselors’ past successful recommendations for college applicants.
3 The Proposed Single-agent CFE Generators
This section outlines the single-agent CFE generators for the proposed hl-continuous and hl-discrete CFEs. Each generator relies on predefined, real-world-like actions to solve optimization problems for CFE generation: the weighted set cover problem for hl-discrete CFEs and integer linear programming for hl-continuous CFEs.
3.1 The Single-agent hl-continuous CFE Generation
Below, we formally define hl-continuous actions and the single-agent hl-continuous CFE generation process.
Definition 1.
(hl-continuous action): An hl-continuous action is a signed ($\pm$) and predefined real-world action whose cost and varied effects on an agent’s input features are predefined and known. For example, in Figure 1, the hl-continuous action: $ca_{1}$: take “leavening agents: cream of tartar” modifies $6+$ features by a known amount and the agent incurs a cost (e.g., estimated average price in USD) that is known apriori.
The singe-agent hl-continuous CFE generator. This generator produces an hl-continuous CFE by solving an integer linear program (ILP). Given the profile of a negatively classified agent $\mathbf{x}$ and a set of hl-continuous actions with known costs (defined above), the objective is to identify the lowest-cost subset of hl-continuous actions that, when taken, modify the agent’s features to achieve a positive classification. The ILP is of the form:
minimize $\sum_{j\in J}\text{cost}_{j}a_{j}$ (2)
s.t. $\mathbf{c}^{T}\sum_{j\in J}a_{j}\cdot(2\epsilon_{j}-1)\cdot\mathbf{v}_{j}\geq-(\mathbf{c}^{T}\mathbf{x}+b)+\delta$
$\epsilon_{j}\in\{0,1\},\quad a_{j}\in\{0,1\},\quad\forall j\in J$
where $J$ denotes the indices of the hl-continuous actions, with each action represented by a vector $\mathbf{v}_{j}$ and with a predefined cost, $\text{cost}_{j}\in\mathbb{R}_{+}$. The boolean variable $a_{j}$ indicates the inclusion ($a_{j}=1$) or exclusion ($a_{j}=0$) of the $j^{\textrm{th}}$ hl-continuous action, while $\epsilon_{j}$ encodes the sign of this action, representing addition ($\epsilon_{j}=1$) or subtraction ($\epsilon_{j}=0$). The coefficients $\mathbf{c}$ and intercept $b$ are the parameters of the linear classifier, and $\delta$ is a small positive value that ensures strict inequality.
3.2 The Single-agent hl-discrete CFE Generation
Below, we formally define hl-discrete actions and the single-agent hl-discrete CFE generation process.